OpenStudy (anonymous):

If an object is accelerating at a rate of 8m/s2, and the force increases by a factor of 4 and the mass increases by a factor of 2, what is the new acceleration?,

OpenStudy (ghazi):

\[F=ma\] now if new F = 4F and new m = 2m then \[A _{n}=\frac{ F _{n} }{ m _{n} }=\frac{ 4F }{ 2m }=2(\frac{ F }{ m })=2a\] An is new acceleration therefore An=2*8=16m/s^2